Problem solving extending charts!
1. Find the number of full 8 hr. shifts that Maria worked last month
(1) She worked twice as many 6 hr. shifts as 8 hr. shifts.
(2) She worked a total of 280 hrs.
Then you make a chart One column is 6h shifts the other is 8 h shifts. The first row is Hrs. Per shift x Number of shifts = Total hrs. worked.

Answer by bam878s(77)

Let x=number of 8hr shifts and y=number of 6hr shifts.
Note y=2x
we have 8x+6y=280
sub for y
8x+6(2x)=280
8x+12x=280
20x=280
divide both sides by 20 to get
x=14 8hr shifts
y=14*2=28 6hr shifts